Short Report – QA/QC Meeting Apr 2019

Meeting Date: 2 May 2019


LH, SG, IF, MG, RM, TB, PM, PL, EPL (9)


Talking Points


  • If you are (Tech-)SRP, please attend QA/QC meetings or tell LH if you can’t, needed for planning of the meetings.
  • SRP & Tech-SRP: please prepare short statement about your site and send to to LH (max. 2 sentences)
  • The purpose of QA/QC meetings is to check on current, incoming data. SRPs choose specific issues we should look at together and discuss in the group. Fluxes are checked if the respective SRP wishes to do so.

Example of shifted timestamps

CH-LAE: Example of shifted timestamps in 2018. These kind of errors can easily be seen in fingerprint / heatmap plots, but not in “normal” time series plots (e.g. scatter plots, line plots).


Site Reports


SRP: –
T-SRP: –


SRP (IF): QCL outages in Jan and Feb followed by a decrease in measurement frequency to 5 Hz, dropping to 1 Hz over Easter. Now this should be fixed, will check tomorrow (3 May). CO2 looks good, uptake started. Technicians took the LI-7500 down and calibrated it yesterday (1 May). QCL measurement frequency problems were solved by end of April. Power test on 30th April responsible for a short data gap in fluxes. Last heat flux plate stopped working at the beginning of April, needs to be replaced.
T-SRP (PL): QCL data had gaps because of connectivity and measurement frequency problems. New SWC and TS profiles were installed on 12 Apr 2019. LI-7500 has been calibrated.


SRP (MG, LH): No mjor issues, EC system running fine. CO2 uptake started in April.
T-SRP (PM): Site was running smoothly. Snow melting period.


SRP (LH): No major issues, strong CO2 uptake started in April.
T-SRP (PL): EC data is complete. G_GF1_0x06_2 needs replacing. IfDews Station has been installed on 03.05.2019


SRP (RM): Large data gap due to malfunction of the laser. Since two weeks everything works again.
T-SRP: –


SRP: –
T-SRP: –


SRP (RM): Some data gaps due to IRGA problems. However, data looks ok.
T-SRP: –


Last Updated on 31 May 2019 17:50